# On Geometry and Topology of Hessian Manifolds

### v4

@@ 4.3 Koszul’s Uncertainty Principle
Let $((M,\nabla),h)$ be an $n$-dimensional Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then the identity
$\operatorname{div}_{h}(\eta^{\sharp})=-\varkappa(\eta^{\sharp})+n$
holds, where $\sharp$ is the musical isomorphism induce by $h$.
@@
Denote by $A$ the Amari–Chentsov tensor of $((M,\nabla),h)$ and denote by
$\bar{\nabla}:=\frac{1}{2}(\nabla+\nabla^{*})$
the Levi-Civita connection of $(M,h)$. Then, by definition, we have
$(\bar{\nabla}_{X}\eta)(Y)=h(X,Y)-\frac{1}{2}A(X,Y,\eta^{\sharp}).$
Therefore, by Lemma 4.21, we conclude that
$\operatorname{div}_{h}(\eta^{\sharp})=-\star d\star\eta=-\varkappa(\eta^{\sharp})+n$
where $\star$ is the Hodge star operator on $(M,h)$. ∎
As usual, for a compact Riemannian manifold $(M,g)$, we denote by $\operatorname{Vol}_{g}(M)$ is volume.
Corollary 4.23.
Let $((M,\nabla),h)$ be an $n$-dimensional compact Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then the integral identity
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta$
holds, where $\star$ is the Hodge star operator on $(M,h)$.
@@
Apply divergence theorem to the formula in Lemma 4.22. ∎
With the preparations done, the geometric uncertainty principle can now be derived easily as a corollary:
Corollary 4.24.
Let $((M,\nabla),h)$ be an $n$-dimensional compact Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then the integral inequality
$\operatorname{Vol}_{h}(M)^{2}\leq\frac{1}{n^{2}}\left(\int_{M}|\varkappa|^{2}\operatorname{vol}_{h}\right)\left(\int_{M}|\eta|^{2}\operatorname{vol}_{h}\right)$
holds on $M$, with equality if and only if $\varkappa=k\cdot\eta$ for some positive real number $k\in\mathbb{R}$, that is, $h$ is a Cheng-Yau metric.
Proof.
Applying Cauchy-Schwarz inequality to the integral formula in Corollary 4.23, we immediately obatin the desired integral inequality. The equality holds exactly when there exists a real constant $k\in\mathbb{R}$ such that $\varkappa=k\cdot\eta$. By Corollary 4.23, we have
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta=\frac{k}{n}\int_{M}\eta\wedge\star\eta=\frac{k}{n}\int_{M}|\eta|^{2}\operatorname{vol}_{h}.$
Therefore the constant $k\in\mathbb{R}$ must be positive. ∎
A facile application of Corollary 4.24 is the following non-vanishing theorem for Riemann curvature tensors of compact Hessian manifolds of Koszul type.
Corollary 4.25.
Compact Hessian manifolds of Koszul type are not Riemannian flat.
Proof.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form.
Assume the contrary that $(M,h)$ is Riemannian flat. Then, by Theorem 2.44, we have that $\varkappa=0$. But Corollary 4.24 then yields $\operatorname{Vol}_{h}(M)\leq 0$, which is absurd. ∎
Remark 4.26.
The integral inequality stated in Corollary 4.24 is an analog of Heisenberg uncertainty principle in quantum mechanics: For a given dimension and volume, a $1$-form potential small in $L^{2}$-norm must be compensated by a first Koszul form large in $L^{2}$-norm, and vice-versa. This phenomenon indicates a highly non-trivial relation between a $1$-form potential and the first Koszul form of a Hessian manifold. We name this integral inequality as Koszul’s uncertainty principle, in honor of the great French mathematician Jean-Louis Koszul.
Remark 4.27.
On a compact Hessian manifold $((M,\nabla),h)$ of Koszul type, the integral inequality of Koszul’s uncertainty principle is sharp, and the equality is achieved exactly when $h$ is a Cheng-Yau metric. This reveals a deep connection with Theorem 4.14 and Proposition 4.19.
5 Topics in Low Dimensional Hessian Geometry
5.1 Exotic Potentials for the Euclidean Plane
Hessian potential functions for a metric, when exist, are in general not unique. For instance, besides the half norm square, the flat metric tensor of Euclidean or Minkowski spaces may well be the Hessian quadratic form of other potential functions.
Here, we shall provide a convenient equivalent condition for a smooth function defined on a planar domain to be a Hessian potential of a flat metric.
Lemma 5.1.
Let $\Omega$ be a domain in $\mathbb{R}^{2}$, and let
$\textup{{x}}\colon\Omega\rightarrow\mathbb{R}^{3},\quad(u,v)\mapsto\textup{{x}}(u,v)$
be a parametrization of a regular smooth surface $S$ in $\mathbb{R}^{3}$. Suppose that $S$ does not pass through the origin, and the vectors $\textup{{x}},\textup{{x}}_{u},\textup{{x}}_{v}$ are everywhere $\mathbb{R}$-linearly dependent. Then, there exists a homogeneous function $F$ in three variables such that $F(\textup{{x}})\equiv 0$. Consequently, $S$ lies on a cone of which apex is the origin.
Proof.
This is a well-known fact in classical differential geometry in $\mathbb{R}^{3}$. We nevertheless give a sketch of proof.
Consider the smooth mapping $\phi:=\textbf{x}/\|\textbf{x}\|$. By the assumption on vectors $\textbf{x},\textbf{x}_{u},\textbf{x}_{v}$, a straightforward computation shows that $\phi_{u}\times\phi_{v}\equiv 0$. Therefore, the image $\phi(\Omega)$ is a smooth curve on the $2$-sphere $\mathbb{S}^{2}:=\{(x,y,z)\in\mathbb{R}^{3}|x^{2}+y^{2}+z^{2}=1\}$. By the implicit function theorem, there exists a smooth function $f\colon\mathbb{S}^{2}\rightarrow\mathbb{R}$ such that $\phi(\Omega)$ is locally the zero locus of $f$. Now, define homogeneous function $F$ of degree zero via
$F(x,y,z):=f(x/r,y/r,z/r)$
where $r=\sqrt{x^{2}+y^{2}+z^{2}}$. Then, by the very construction, we have $F(\textbf{x})\equiv 0$. ∎
Proposition 5.2.
Let $f$ be a smooth function defined on a domain in $\mathbb{R}^{2}$ such that $\operatorname{Hess}(f)$ is everywhere non-degenerate, so in particular
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a pseudo-Riemannian metric. Then, $g$ is flat if and only if there exists a homogeneous function $F$ in three variables such that $F(f_{xx},f_{xy},f_{yy})\equiv 0$.
Proof.
Combine Corollary 4.4 with Lemma 5.1. ∎
Example 5.3.
Let $f$ be a smooth homogeneous function of degree $d$ defined on a domain in $\mathbb{R}^{2}$ such that its Hessian quadratic form $H:=\operatorname{Hess}(f)$ is everywhere non-degenerate. Then, by Euler’s theorem on homogeneous functions, we have
$(2-d)H+x\frac{\partial H}{\partial x}+y\frac{\partial H}{\partial y}=0$
that is, the matrices $H,H_{x},H_{y}$ are everywhere $\mathbb{R}$-linearly dependent. Therefore, by Corollary 4.4, we have that
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a flat pseudo-Riemannian metric.
The following interesting example shows that nonlinear harmonic functions are Hessian potentials for the flat $1+1$ Minkowski plane $\mathbb{R}^{1,1}$.
Example 5.4.
Let $\Omega$ be a domain in $\mathbb{C}^{1}$, and let $f\in C^{\infty}(\Omega)$ be a nonlinear harmonic function. Then, $\operatorname{tr}(\operatorname{Hess}(f))=f_{xx}+f_{yy}=0$ and
$\operatorname{det}(\operatorname{Hess}(f))=f_{xx}f_{yy}-f_{xy}^{2}=-f_{yy}^{2}-f_{xy}^{2}<0.$
By Cayley–Hamilton theorem, we obtain that $\operatorname{Hess}(f)$ is non-degenerate and indefinite, and hence
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a Lorentzian metric. Moreover, by Proposition 5.2, $g$ is flat, i.e. locally the metric of the $1+1$ Minkowski space.
We also suggest the following very concrete example of a Hessian potential for the flat Euclidean plane, which is constructed using the method introduced in Theorem 1.8.
Example 5.5.
Let $\mathbb{H}^{2}:=\{(x,y)\in\mathbb{R}^{2}:y>0\}$ be the upper half plane and define $f\in C^{\infty}(\mathbb{H}^{2})$ by
$f(x,y):=\frac{x^{2}}{2y}+\frac{1}{4}\operatorname{log}(y)y.$
Then the coordinate change $(x(r,\theta),y(r,\theta))=(r^{2}\theta,r^{2})$ brings the Hessian quadratic form
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}=\frac{1}{y}dx^{2}-\frac{2x}{y^{2}}dxdy+\frac{4x^{2}+y^{2}}{4y^{3}}dy^{2}$
of $f$ to the flat metric tensor $dr^{2}+r^{2}d\theta^{2}$ of the Euclidean plane in its polar coordinates. The flatness of $g$ can be proved also by noticing that
$f_{xx}^{2}-4f_{xx}f_{yy}+4f_{xy}^{2}\equiv 0$
and then apply Proposition 5.2.
Recall from Example 5.3 that, in dimension $2$, all smooth homogeneous functions with non-singular Hessian are potentials for flat metrics. However, a computation in [30] indicates that this is no longer true in higher dimension.
Before we close this section, we cite from [30] an example given by Dubrovin, to show that techniques from algebraic geometry, for example the Clebsch covariant of projective plane curves, might be useful for finding homogeneous potentials.
Example 5.6.
Let $f(x,y,z):=x^{6}+y^{6}+z^{6}-10(x^{3}y^{3}+y^{3}z^{3}+z^{3}x^{3})$ be the Maschke sextic. Then the Hessian quadratic form
$g=\sum_{i=1}^{3}\sum_{j=1}^{3}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}dx_{i}\otimes dx_{j}$
of $f$ is a flat metric on $\mathbb{R}^{3}$.
For more discussion on homogeneous polynomial potential of flat spaces, we refer to the work of B.Totaro [30].
@@ 5.2 Complete Hessian Surfaces
Let $(M,\nabla)$ be an affine manifold. We denote by $\operatorname{Aff}(M;\nabla)$ the group of affine automorphisms of $(M,\nabla)$. Whenever $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$, and $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space, by convention we omit $\nabla$ to write $\operatorname{Aff}(\Omega):=\operatorname{Aff}(\Omega;\nabla)$.
@@
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$ and $\operatorname{recc}(\Omega)$ its recession cone. Let $f\in\operatorname{Aff}_{2}(\mathbb{R})$ and let $T$ be the linear part of $f$, that is, $f(\textup{{x}})=T\textup{{x}}+f(0)$ for all $\textup{{x}}\in\mathbb{R}^{d}$. Suppose that $\Omega=f(\Omega)$. Then $T\textup{{x}}\in\operatorname{recc}(\Omega)$ for all $\textup{{x}}\in\operatorname{recc}(\Omega)$.
@@
Take any $\textbf{x}\in\operatorname{recc}(\Omega)$ and $\textbf{y}\in\Omega$. By the definition of a recession cone, it suffices to prove that $T\textbf{x}+\textbf{y}\in\Omega$. But indeed, we have
$T\textbf{x}+\textbf{y}=T\textbf{x}+Tf^{-1}(\textbf{y})+f(0)=f(\textbf{x}+f^{-1}(\textbf{y}))\in\Omega$
@@
The following proposition concerning compactness of the affine automorphism group of a bounded convex domain is itself interesting and useful.
@@
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to one of the following smooth manifolds:
@@
Since $((M,\nabla),h)$ is a complete Hessian manifold of dimension 2, by Theorem 2.13, there exists a planar convex domain $\Omega$ and a discrete subgroup $G$ of the affine group $\operatorname{Aff}_{2}(\mathbb{R})$, such that $G$ acts freely and proper discontinuously on $\Omega$ by affine transform, and that $M\cong\Omega/G$.
If $G$ is trivial, then $M=\Omega$ is homeomorphic to $\mathbb{R}^{2}$. If $M$ is closed, then by Theorem 2.17, we have that $\chi(M)=0$ and hence $M$ is homeomorphic to the 2-torus or the Klein bottle. Recall from low-dimensional topology that the fundamental group of a non-compact surface is always free. From now on, we assume that $M$ is non-compact and $G$ is a non-trivial free group.
@@
Assume that $\operatorname{recc}(\Omega)$ is a point. Then, by convex geometry, we have that $\Omega$ is bounded. Therefore, by Proposition 5.9, $\operatorname{Aff}(\Omega)$ is a compact Lie group. But, by assumption, $G$ is a infinite discrete subgroup of $\operatorname{Aff}(\Omega)$, which leads to a contradiction. Therefore $\operatorname{recc}(\Omega)$ cannot be a point.
@@
Example 5.12.
A complete Hessian surface of constant Gaußian curvature 1-1 can be isometric to a hyperbolic funnel, that is, the quotient of the Lobachevsky plane by an expanding homothety $z\mapsto\lambda\cdot z$ where $\lambda>1$.
In the polar coordinates $(r,\theta)$ of the complex plane, the $1$-form potential of a hyperbolic funnel is precisely
$\eta=\frac{dr}{2r}+\left(\frac{1}{2}\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)}-\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)^{3}}\right)d\theta$
defined on an open annulus centered at the origin. A straightforward computation shows that $\eta$ is closed, but not exact.
5.3 Compact Orientable Hessian 3-folds
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In is section, we construct concretely a family of compact radiant Hessian manifolds of Koszul type with negative scalar curvature.

### v6

@@ 4.3 Convex Decomposition Lemma
Let $((M,D),g)$ be a Hessian manifold and $(N,\nabla)$ an affine manifold. Suppose that $i\colon N\rightarrow M$ is an affine immersion. Then $((N,\nabla),i^{*}g)$ is a Hessian manifold.
@@
Since $i\colon N\rightarrow M$ is affine, we have that $\nabla=i^{*}D$, and hence $\nabla(i^{*}g)=i^{*}(Dg)$. Since $((M,D),g)$ is a Hessian manifold, we obtain that $\nabla(i^{*}g)$ is totally symmetric by Theorem 1.8. Therefore $\nabla(i^{*}g)$ is totally symmetric as $Dg$ is. Again by Theorem 1.8, we conclude that $((N,\nabla),i^{*}g)$ is a Hessian manifold. ∎
We will also need the following result on the famous Auslander conjecture.
Theorem 4.14 (Abels et al.,2002[1]).
Let $(M,\nabla)$ be a compact affine manifold of dimension $n\leq 6$, and $\pi\colon\tilde{M}\rightarrow M$ the universal covering of $M$. Let $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ be the affine development of $(\tilde{M},\pi^{*}\nabla)$. Suppose that $\Phi$ is a diffeomorphism. Then $(M,\nabla)$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Recall that every convex domain in $\mathbb{R}^{d}$ can be written as the Cartesian product of a regular convex domain and a affine subspace of $\mathbb{R}^{d}$. This fact motivates the following splitting theorem for compact Hessian manifolds.
Theorem 4.15.
[…]
$((E,f^{*}\nabla),f^{*}h)=((B\times\mathbb{T}^{n-k},\hat{\nabla}\oplus D),g+ds^{2})$
[…]
@@
[…]
Since $\Omega$ contains no affine subspaces of positive dimension, for any $\gamma\in\pi_{1}(M)$, there exists $\phi_{\gamma}\in\operatorname{Aff}(\Omega)$ and $U_{\gamma}\in\operatorname{GL}(\mathbb{R}^{d})$ such that
$\gamma(\textbf{x},\textbf{y})=(\phi_{\gamma}(\textbf{x}),T_{\gamma}\textbf{x}+U_{\gamma}\textbf{y}+\textbf{v}_{\gamma})$
for some real $(n-d)\times d$-matrix $T_{\gamma}$ and some vector $\textbf{v}_{\gamma}\in\mathbb{R}^{d}$. Consider the homomorphism
$\phi\colon\pi_{1}(M)\rightarrow\operatorname{Aff}(\Omega),\quad\gamma\mapsto\phi_{\gamma}.$
For each $\textbf{x}\in\Omega$ denote
$N_{\textbf{x}}:=\frac{\{\textbf{x}\}\times\mathbb{R}^{d}}{\operatorname{Ker}(\phi)}$
where the normal subgroup $K:=\operatorname{Ker}(\phi)$ of $\pi_{1}(M)$ acts naturally on the vertical fiber $\{\textbf{x}\}\times\mathbb{R}^{d}$ by affine transformation.
[…]
Therefore, by Theorem 2.4, there exists a convex function $\varphi\in C^{\infty}(\Omega\times\mathbb{R}^{d})$ such that
$\pi^{*}h=\sum_{i=1}^{n-d}\sum_{j=1}^{n-d}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}dx_{i}dx_{j}+2\sum_{i=1}^{n-d}\sum_{j=1}^{d}\frac{\partial^{2}\varphi}{\partial x_{i}\partial y_{j}}dx_{i}dy_{j}+\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}dy_{i}dy_{j}$
and moreover,
$(\pi^{*}h)|_{\{\textbf{a}\}\times\mathbb{R}^{d}}=\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}(a_{1},\dots,a_{n-d},y_{1},\dots,y_{d})dy_{i}\otimes dy_{j}$
is a flat Riemannian metric for any $\textbf{a}=(a_{1},\dots,a_{n-d})\in\Omega$. Therefore, by Corollary 1.11, we obtain that
$\frac{\partial^{3}\varphi}{\partial y_{i}\partial y_{j}\partial y_{k}}=0$
for all $i,j,k\in\{1,\dots,d\}$, and hence there exist a smooth function $u\in C^{\infty}(\Omega)$, a vector-valued function $\xi\in C^{\infty}(\Omega)^{\oplus d}$, and a matrix-valued function $F\in C^{\infty}(\Omega)^{d\times d}$ such that $\varphi(\textbf{x},\textbf{y})=u(\textbf{x})+\langle\xi(\textbf{x}),\textbf{y}\rangle+\langle F(\textbf{x})\textbf{y},\textbf{y}\rangle$ for all $(\textbf{x},\textbf{y})\in\Omega\times\mathbb{R}^{d}$.
Recall that an element $\gamma\in K$ acts on $\Omega\times\mathbb{R}^{d}$ by pure translation $\gamma(\textbf{x},\textbf{y})=(\textbf{x},T_{\gamma}\textbf{x}+\textbf{y}+\textbf{v}_{\gamma})$. The $K$-invariance of the Hessian quadratic form of $\varphi$ then yields conservation laws
$\frac{\partial}{\partial x_{i}}\frac{\partial}{\partial x_{j}}(F(\textbf{x})\xi(\textbf{x}))=\frac{\partial}{\partial x_{k}}\left(F(\textbf{x})(T_{\gamma}\textbf{x}+\textbf{v}_{\gamma})\right)=0$
for all $\gamma\in K$ and all $i,j,k\in\{1,\dots,n-d\}$, which implies that there exist real $d\times d$-matrices $A_{0},\dots,A_{n-d}$ and an affine mapping $\phi\colon\mathbb{R}^{n-d}\rightarrow\mathbb{R}^{d}$ such that $F(\textbf{x})=L(\textbf{x})^{-1}$ and $\xi(\textbf{x})=L(\textbf{x})^{-1}\phi(\textbf{x})$ where
$L(x_{1},\dots,x_{n-d}):=A_{0}+\sum_{i=1}^{n-d}x_{i}A_{i}$
and hence, up to an affine change of coordinates, the potential $\varphi$ is of the form
$\varphi(\textbf{x},\textbf{y}):=u(\textbf{x})+\langle L(\textbf{x})^{-1}\textbf{y},\textbf{y}\rangle\in C^{\infty}(\Omega)\otimes\mathbb{R}[y_{1},\cdots,y_{d}].$
In particular, for any $\textbf{x}\in\Omega$, the matrix $L(\textbf{x})$ is symmetric and positive definite.
Take any arbitrary $\gamma\in\pi_{1}(M)$. Since the Hessian quadratic form of $\varphi\circ\gamma-\varphi$ vanishes, we have that $\varphi\circ\gamma-\varphi$ is an affine function. Now, we regard $\varphi(\textbf{x},\textbf{y})\in(C^{\infty}(\Omega))[y_{1},\cdots,y_{d}]$ as a quadratic polynomial. The quadratic term of $\varphi$ and $\varphi\circ\gamma$ then coincide, that is,
$L(S_{\gamma}\textbf{x})=U_{\gamma}^{\dagger}L(\textbf{x})U_{\gamma}$
for all $\textbf{x}\in\Omega$, where $U_{\gamma}^{\dagger}$ here is the transpose of $U_{\gamma}$. By considering the limit $\textbf{x}\rightarrow 0$, we have
$A_{0}=L(0)=U_{\gamma}^{\dagger}L(0)U_{\gamma}=U_{\gamma}^{\dagger}A_{0}U_{\gamma}$
and $A_{0}=L(0)\in\operatorname{GL}(\mathbb{R}^{d})$ is a positive definite symmetric matrix. Therefore, we conclude that
$U_{\gamma}\in\operatorname{SL}_{d}(\mathbb{Z})\cap\operatorname{O}(d).$
Since $\operatorname{SL}_{d}(\mathbb{Z})$ is discrete and $\operatorname{O}(d)$ is compact, we have that $\operatorname{SL}_{d}(\mathbb{Z})\cap\operatorname{O}(d)$ is a finite group. By lifting to a finite-sheet covering space of $M$, we may assume that $U_{\gamma}$ is the identity matrix for all $\gamma\in\pi_{1}(M)$, and hence $L(\textbf{x})=L(S_{\gamma}\textbf{x})$ for all $\textbf{x}\in\Omega$.
Now, denote $\textbf{v}:=(\operatorname{tr}(A_{1}),\dots,\operatorname{tr}(A_{n-d}))\in\mathbb{R}^{n-d}$. Then
$\operatorname{tr}(L(\textbf{x}))=\langle\textbf{x},\textbf{v}\rangle+\operatorname{tr}(A_{0}),$
and $L(\textbf{x})=L(S_{\gamma}\textbf{x})$ implies that $\langle\textbf{x},\textbf{v}\rangle=\langle S_{\gamma}\textbf{x},\textbf{v}\rangle$ for all $\textbf{x}\in\Omega$. Therefore, $C^{\infty}(\Omega)^{H}$ is $H$-invariant. Since the $H$-action on $\Omega$ is cocompact, we have that the affine function $L(\textbf{x})$ is bounded, and hence must be constant. Therefore, we obtain $\langle\textbf{x},\textbf{v}\rangle=0$ for all $\textbf{x}\in\Omega$ and hence $\textbf{v}=0$, i.e. $\operatorname{tr}(A_{i})=0$ for all $i=1,\dots,n-d$.
Take any $\textbf{x}=(x_{1},\dots,x_{n-d})\in\Omega$. If
$A(x_{1},\dots,x_{n-d}):=\sum_{i=1}^{n-d}x_{i}A_{i}$
admits a negative eigenvalue, then for sufficient large $t>0$, we have
$L(\textbf{x}\cdot t)/t=A(\textbf{x})+\frac{1}{t}A_{0}$
is not positive definite, contradiction.
Therefore, for any $\textbf{x}=(x_{1},\dots,x_{n-d})\in\Omega$, we have
$A(\textbf{x})=\sum_{i=1}^{n-d}x_{i}A_{i}\geq 0.$
But, since $\operatorname{tr}(A_{1})=\cdots=\operatorname{tr}(A_{n-d})=0$, we have $\operatorname{tr}(A(\textbf{x}))=0$ and hence $A(\textbf{x})=0$ for all $\textbf{x}\in\Omega$. Therefore, we conclude that $L(\textbf{x})\equiv A_{0}$ is a constant matrix-valued function, and hence, up to an affine change of coordinates, we may assume that
$\varphi(\textbf{x},\textbf{y})=u(\textbf{x})+\frac{1}{2}\|\textbf{y}\|^{2}.$
The equation
$\frac{\partial\,T_{\gamma}\textbf{x}}{\partial x_{i}}=\frac{1}{2}\frac{\partial}{\partial x_{i}}\left(F(\textbf{x})(T_{\gamma}\textbf{x}+\textbf{v}_{\gamma})\right)=0,$
where $i\in\{1,\dots,n-d\}$ and $\gamma\in\pi_{1}(M)$, now gives that $T_{\gamma}=0$ for all $\gamma\in\pi_{1}(M)$. In particular, the central extension
$1\longrightarrow K\longrightarrow\pi_{1}(M)\longrightarrow H\longrightarrow 1$
splits as a direct product $\pi_{1}(M)=H\times K$. The theorem is therefore proved. ∎
5 Classification Results in Low Dimensions
@@ 5.1 Complete Hessian Surfaces
For a convex domain $\Omega$ in $\mathbb{R}^{d}$, we denote by
$\operatorname{Aff}(\Omega):=\{f\in\operatorname{Aff}_{d}(\mathbb{R}):f(\Omega)=\Omega\}$
the group of affine automorphisms of $\Omega$.
@@
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$ and $\operatorname{recc}(\Omega)$ its recession cone. Let $f\in\operatorname{Aff}_{2}(\mathbb{R})$ and let $A$ be the linear part of $f$, that is, $f(\textup{{x}})=A\textup{{x}}+f(0)$ for all $\textup{{x}}\in\mathbb{R}^{d}$. Suppose that $\Omega=f(\Omega)$. Then $A\textup{{x}}\in\operatorname{recc}(\Omega)$ for all $\textup{{x}}\in\operatorname{recc}(\Omega)$.
@@
Take any $\textbf{x}\in\operatorname{recc}(\Omega)$ and $\textbf{y}\in\Omega$. By the definition of a recession cone, it suffices to prove that $A\textbf{x}+\textbf{y}\in\Omega$. But indeed, we have
$A\textbf{x}+\textbf{y}=A\textbf{x}+Af^{-1}(\textbf{y})+f(0)=f(\textbf{x}+f^{-1}(\textbf{y}))\in\Omega$
@@
The following proposition concerning compactness of the affine automorphism group of a convex body is itself interesting and useful.
@@
Lemma 5.4.
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{d}$ and let a subgroup $G$ of $\operatorname{Aff}(\Omega)$ act on $\Omega$ by affine transformation. Then the quotient space $\Omega/G$ is never compact.
Proof.
By Proposition 5.3, up to an affine change of coordinates, we may assume $G$ is a subgroup of the orthogonal group $\operatorname{O}(d)$, and that $0\in\mathbb{R}^{d}$ is the center of mass of $\Omega$. Denote by $\partial\Omega$ the boundary of $\Omega$ in $\mathbb{R}^{d}$. Then the $G$-action on $\Omega$ extends uniquely to $\partial\Omega$. Notice that
$\frac{\partial\Omega\times[0,1)}{\partial\Omega\times\{0\}}\cong\{t\textbf{x}\in\mathbb{R}^{d}:\textbf{x}\in\partial\Omega,0\leq t<1\}=\Omega.$
Then we conclude that
$\Omega/G\cong\frac{(\partial\Omega/G)\times[0,1)}{(\partial\Omega/G)\times\{0\}}$
is non-compact, as it admits a continuous surjective mapping onto the interval $[0,1)$. ∎
@@
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to a complete flat Riemannian manifold of dimension two, that is, one of the following smooth manifolds:
@@
Since $((M,\nabla),h)$ is a complete Hessian manifold of dimension 2, by Theorem 2.12, there exists a planar convex domain $\Omega$ and a discrete subgroup $G$ of the affine group $\operatorname{Aff}_{2}(\mathbb{R})$, such that $G$ acts freely and proper discontinuously on $\Omega$ by affine transform so that $M\cong\Omega/G$.
If $G$ is trivial, then $M=\Omega$ is homeomorphic to $\mathbb{R}^{2}$. If $M$ is closed, then by Proposition 4.3, we have that $\chi(M)=0$ and hence $M$ is homeomorphic to the 2-torus or the Klein bottle. Recall from low-dimensional topology that the fundamental group of a non-compact surface is always free. From now on, we assume that $M$ is non-compact and $G$ is a non-trivial free group.
@@
Recall from convex geometry that a convex domain is bounded if and only if its recession cone is a point. If $\operatorname{recc}(\Omega)$ is indeed a point then the boundedness of $\Omega$ violates Lemma 5.4. Therefore $\operatorname{recc}(\Omega)$ cannot be a point.
@@
5.2 Closed Orientable Hessian 3-Manifolds
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In this section, we construct concretely a family of compact radiant Hessian manifolds of Koszul type with negative scalar curvature.